OpenStudy (anonymous):

Give the radius and center for the following equation. 3x^2+3y^2=75 center: (____,____) radius=_____

OpenStudy (anonymous):

du try with ,or what u get,?

OpenStudy (anonymous):

I did n`t get that...

OpenStudy (ash2326):

Equation of the form \[(x-a)^2+(y-b)^2=r^2\] represents a circle with center (a, b) and radius r We are given \[3x^2+3y^2=75\] First step is to divide both sides by 3 \[x^2+y^2=25\] Now @super_map Could you try to find the center and radius of the circle?

OpenStudy (anonymous):

(0,0) r=5

OpenStudy (ash2326):

@badi Please don't just give answers, explain how do you get the answer Thanks:)

OpenStudy (anonymous):

thank you :)

OpenStudy (ash2326):

@super_map did you understand?

OpenStudy (anonymous):

yes I did, thank you very much!

OpenStudy (ash2326):

You're welcome:D